Use bond enthalpies (in kJ/mol): C-C 348 C=C 614 C≡C 839 H-H 436 C-H 413 O-H 463 C-O 358 C=O 799 to estimate ΔH for the reaction H2C=CH–CH2–OH (g) = > HC≡C–CH=O(g) + 2 H2(g)
Answer #1

4 = CH=CH-он ис =C-CH=0 + ан, ДА = 2( 26 ) + 3 (13) +749 +8 39 + ч 8 – s (чig) - віч - 64 - 348 - 3sg ДН 2 Чs C6 - 3848 ОН =